

********************************************************************
UARS WEEKLY STATUS REPORT - 06 March 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data during this past week without apparent problems.
SUSIM dwelled on the strong Lyman alpha emission line during last
Thursday's solar eclipse which, at maximum, was about 85% of totality.
***********
PEM
***********
PEM/AXIS was activated March 2 and is now producing energy deposition
profiles from electrons inverted from the AXIS X-ray measurements.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
***********
HALOE
***********
HALOE was powered on March 4, 1998. All engineering parameters are nominal
and the science channel outputs are within expected values.
***********
MLS
***********
The MLS instrument has been powered off since Mar 3. MLS is scheduled to
resume south viewing operations on Mar 12.
***********
HRDI
***********
HRDI is currently off. Next expected turn-on is March 7.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
As part of the UARS re-engineering effort, two CISCO routers
are being installed with the MPG RAC. These routers are being
set up so that the Command Management System and the MPG RAC
will no longer depend on the routers located in Building 23
with the UARS CDHF.
***********
CDHF
***********
The CDHF has received a total of 376 CD-ROM requests as of 06-March-1998.
Four orders were received and shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) PROD (1379R-979C=400L)
123 days MLS (4.90) PROD (1297R-804C=493L)
8 days CLAES (7.0) ENG (17R-9C=8L)
The CDHF continues to work with NSI to develop a support plan for the
entire UARSNET network, focusing on Instrument RACs first, redesign issues,
and then the global user community.
The CDHF will load January 1998 current day DAAC data in March 1998.
Pending software upgrades: IDL 5.0, OpenIngres, UCSS.
***********
FOT
***********
REPORTING PERIOD
This report covers 21 February 1998 (Orbit 35226, GMT Day 052) through 27
February 1998 (Orbit 35329, GMT Day 058).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
MLS 053/2329 35255 Off
WINDII 057/1027 35307 On
WINDII 057/2004 35313 Off
The beta angle reached a minimum angle of 0.0 degrees on 25 February 1998
(DOY 056). The beta angle is now increasing toward a maximum angle of 54.5
degrees on 15 March 1998 (DOY 074).
Four (4) SSPP SOLSTICE offsets requested by SUSIM operations were performed
this week. One offset took place each day, beginning with GMT Day 052 and
ending with GMT Day 055.
A Reverse-to-Forward Yaw Around maneuver (number 66) was successfully
executed in Orbit 35294 at 14:35:50 UTC on 25 February 1998. The SSPP
Instruments (ACRIM, SOLSTICE, and SUSIM) were enabled. Due to the SA
position at 269 degrees, the effective night length of the spacecraft after
the maneuver was 48 minutes. Post maneuver analysis of data indicate that
during the Yaw Around, the end-of-night load bus voltage was 25.0 volts and
the average depths of discharge (DODs) on Batteries 2 and 3 were
approximately 22.5 and 22.0 percent respectively.
UARS experienced a Solar Eclipse on 26 February 1998 in Orbit 35311 during
battery day at approximately 057/17:07 UTC and exited at approximately
057/17:24 UTC. The maximum sun coverage was 86.3 percent. The WINDII
instrument was powered on four orbits prior to the eclipse to view eclipse-
related data, and powered back off two orbits following the eclipse. ACRIM,
SOLSTICE, and SUSIM were also powered during the eclipse. PMON heater
cycling was disabled during the eclipse to conserve power. Post-eclipse
analysis of data indicate that during the eclipse orbit, approximately 19%
of nominal solar array input energy was lost. During the spacecraft night
following the eclipse, the load bus voltage reached a minimum of 23.7
volts. Power and thermal subsystem stability were restored within one
orbit following the eclipse.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 25.4 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.4 degrees C. Additional battery performance data for
each battery on 21 February 1998 and 27 February 1998 are provided in
Appendix D.
The clock error ranged from -0.7 to +10.7 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 15
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% HRDI 05
% MLS 01
% WINDII 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT none
TRANSPONDER FREQUENCIES:
Transponder A: 98/041 - 2287.497104
Transponder B: 98/042 - 2287.499330
GROUND SYSTEM OPERATION
No (0) operational support problem(s) occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/057 show the observatory data loss to be
20 hours, 42 minutes, 7.540 seconds (no change since last reported). This
is a 0.0367 percent data loss which equals a 99.9633 percent data capture
for the mission.
FOT is investigating the feasibility of participating in the IP Forward
Link testing. Due to the possible risk of the IP conversion devices
corrupting commands, the FOT would like to test the IP link with the UARS
Test and Training Simulator (UTTS) before transmitting any commands to the
spacecraft. The methodology and requirements for performing a command
loop-back test is currently under investigation.
Testing of the latest AP software (modified to support a reduced ATC load)
was attempted this week. A reduced ATC memory size is required to support
the new PMON software. Testing failed, as a command link between MSOCC and
the UTTS could not be established. A second day of testing also failed due
to the same problem. Efforts by ATSC personnel to isolate the source of
the problem proved unsuccessful. Testing will resume next week.
Operational processing of level-0 telemetry analysis files continues in
parallel on both the UMPGAL (new MPG Alpha machine) and CDHF computer
systems. Level-0 file transfer failures (between CDHF and UMPGAL) occurred
on 24 February at about 02:00 UTC and on 27 February at about 19:00 UTC.
File transfers, when they do not fail, are occurring at about 43 kbps
effective rate. This means a production set of level-0 files will
typically take about 4-1/2 hours to transfer from CDHF to UMPGAL, hardly an
acceptable transfer rate for operational products. Until a more timely
link for transfer of level-0 files can be established, the CDHF system will
be considered the operational system for FOT analysis of telemetry data.
OTHER SIGNIFICANT EVENTS
Work on the upgrade of the OBC Power Monitor (PMON) software continued. The
code is being debugged. The PMON telemetry processing subroutine had to
be redesigned due to excessive code usage. The redesign of the telemetry
processing subroutine is table-driven and very efficient. This re-design
could potentially be used to process all OBC telemetry from within INTER
much more efficiently than the existing code, and all from one centralized
piece of code. This approach could free up tons of memory, and allow
greater flexibility and control over OBC telemetry at the same time and,
therefore, should be seriously considered even though an OBC reload would
be required.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
Attended a meeting on 23 February 1998, to discuss the R&D effort to
develop a command system for the TMOC ground system. Collected and
presented current commanding system documentation to Lockheed Martin and
ICS (Interface & Control Systems) personnel.
A meeting was held with Flight Dynamics Facility (FDF) personnel on 25
February 1998 to obtain the results of the current drag predictions
utilizing revised solar cycle 23 predictions, revised Spacecraft weight,
and revised Spacecraft drag area. The new calculations predict that the
585 -1.25 = 583.75 km lower operational limit will be reached as early as
12 July 1998 (-2 sigma) and 03 December 1998 (+2 sigma). The UARS Program
Manger requested that an orbit adjust be executed by August 1998 to assure
that skilled manpower and hardware resources are still available. The FOT
is responsible (by 11 March 1998) for presenting an activity plan and
timeline for accomplishing this maneuver by the planned date. Extensive
ground testing will be required prior to the orbit adjust, accounting for
most of the preparation time for this maneuver.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 02 April 1998.
UARS Flight Operations plans to attend the Program Experience portion of
the Space Power Workshop at JPL in April 1998. UARS FOT plans on
presenting an overview of the UARS batteries and their performance.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 05 August 1998.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 27 February 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
35249 053/1344 98-061 HRDI HTR cycle from 25 to 200 amp-min
35255 053/2329 98-062 MLS Power OFF
35280 055/1532 98-063 HRDI HTR Cycle from 200 to 150 amp-min
35299 056/2146 98-064 OBC/PMON RTS-4 & RTS-70 for V/T Switch
35307 057/1027 98-065 WINDII Power ON
35308 057/1204 98-066 HRDI HTR cycle form 150 to 100 amp-min
35310 057/1530 98-067 OBC/PMON NOOP RTS 04
35313 057/2004 98-068 WINDII Power OFF
35313 057/2004 98-069 HRDI HTR cycle from 100 to 200 amp-min
35315 057/2334 98-070 HRDI HTR cycle from 200 to 300 amp-min
35321 058/0622 98-071 OBC Clk Adj fr 0.00 to 20.6 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 27 February 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ -----
HRIFESURVTMP Y-HI HRDI Survival temperature 35227
HRIFESURVTMP Y-HI HRDI Survival temperature 35246
HRPSEQBIMON Y-HI HRDI Quit Bus monitor 35246,90
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 35246,82,84,
35307-315, 321-29
WIEUTEMP Y-LO WINDII EU Temperature 35277-301-07,16-29
NBRECBI Y-HI NBTR B REC CURRENT 35308,12
WIPBVOLTAGE Y-LO WINDII PB VOLTAGE 35308,12
SOSTMPRNTRN Y-LO ENTRANCE TEMP 35312-14,21
ACTEMPANA2 Y-LO Temp Sensor Electronics 2 35312-13
QTSSPPMX Y-HI SSPP -X Thermostat 35312
HRIFESURVTEMP Y-HI IFE SURVIVAL TEMP 35312-13
HRRADTEMP Y-HI IFE RADIATOR TEMP 35312
HRPSEQBIMON Y-HI QB CURRENT MONITOR 35312
CDUNRG28 Y-LO +28V UNREG. BUS 35312
ACTEMPANA1 Y-LO Temp Sensor Electronics 1 35315
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
none this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 27 February 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
35321 058/06:29 (27 Feb 98) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 27 February 1998
21 Feb 1998 - GMT Day 052 Beta = 16.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-3.99 +0.0/ +0.0 18.6 0.000 0.0
2 +33.6/ -33.6 +7.41/+5.47 +32.4/-15.2 24.8 1.023 23.1
3 +33.6/ -22.4 +5.79/+4.51 +31.6/-16.0 24.8 1.023 22.4
27 Feb 1998 - GMT Day 058 Beta = 9.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/ +0.0 18.6 0.000 0.0
2 +89.6/ 0.0 +5.47/+4.19 +31.6/-13.6 25.8 1.023 21.9
3 +56.0/ -11.2 +4.19/+2.60 +30.4/-15.2 25.4 1.021 21.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 27 February 1998
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-001 FOT Four January/98 14170 Generic Late Acq.
98-004 FOT 34874 028/1332 TAC loss. Op err.
98-005 FOT 34912 032/0147 TAC Crash.
98-006 FOT 34962 034/0949 TAC Crash.
98-008 FOT 34972 035/0156 TAC Crash
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period